An experienced accountant can prepare a tax return in 16 hours. A novice account can do the job in 23 hours. How long will it take them to do the job working together?
A: 1/368 hr
B: 1/39 hr
C: 368/7 hr
D: 368/39 hr

Found 2 solutions by FrankM, Alan3354:
Answer by FrankM(1040)

Pro can do 1/16 return per hour
Novice can do 1/23 return per hour.
Together, 1/16+1/23 = .0625 + .0435 = .10598 returns per hour or the reciprocal, 9.436 hours for one return.
This is choice D.
With no math, answers A,B, and C make no sense. If both were fast, it would be 8 hours, both slow, 11.5 hours. D is the only answer that is between these two extremes.

Answer by Alan3354(69443)

An experienced accountant can prepare a tax return in 16 hours. A novice account can do the job in 23 hours. How long will it take them to do the job working together?
-----------------
A shortcut for work, flow, resistance, and other things "in parallel" is product over sum
----
16*23/(16+23) = 368/39